OpenStudy (anonymous):

The formula for finding the area of a square that has a side length, s, is A = s2. If a square has an area of 40 square units, what is the length of a side?

OpenStudy (mathstudent55):

\(A = s^2\) \(40 = s^2\) \(s^2 = 40\) Since you want s, take the square root of both sides: \(s = \sqrt {40} \) Now simplify the root.

OpenStudy (anonymous):

4 * 10 or is it 2 * 20

OpenStudy (anonymous):

\[10\sqrt{2}\]

OpenStudy (anonymous):

is that right ya think? @pooja195

pooja195 (pooja195):

You have it flipped

OpenStudy (anonymous):

\[2\sqrt{10}\]

pooja195 (pooja195):

There you go :)

OpenStudy (anonymous):

thanks :)))